# Bridge Sampling Diagnostics

### v1

@@ Abstract
In Bayesian statistics, the marginal likelihood is used for model selection and averaging, yet it is often challenging to compute accurately for complex models. Approaches such as bridge sampling, while effective, may suffer from issues of high variability of the estimates. We present how to estimate Monte Carlo standard error (MCSE) for bridge sampling, and how to diagnose the reliability of MCSE estimates using Pareto-$\hat{k}$ and block reshuffling diagnostics without the need to repeatedly re-run full posterior inference. We demonstrate the behavior with increasingly more difficult simulated posteriors and many real posteriors from the posteriordb database.
@@ 1 Introduction
The marginal likelihood is commonly used in Bayesian statistics. The purpose of this paper is not to advocate the use of marginal likelihood for model selection, but as marginal likelihood is used anyway, the goal is to help reducing the amount of computation used with practical diagnostics. Let $\mathcal{M}$ represent the Bayesian model of interest. The marginal likelihood given the observed data $\mathbf{y}$ is
$p(\mathbf{y}\mid\mathcal{M})=\int_{\Theta}p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})\,d\theta,$ (1)
where $p(\mathbf{y}\mid\theta,\mathcal{M})=\prod_{i=1}^{n}p(y_{i}\mid\theta,\mathcal{M})$ is the likelihood of $\theta$ given observed data $y_{i}$ and the model $\mathcal{M}$, $p(\theta\mid\mathcal{M})$ is the prior distribution of the parameters $\theta$, and $\Theta$ is the parameter space. For all but the simplest cases, this integral becomes intractable, requiring numerical methods to approximate it.
We examine the finite sample size accuracy of bridge sampling algorithm for estimating the marginal likelihood $p(\mathbf{y}\mid\mathcal{M})$. We (i) derive a closed-form expression for the Monte-Carlo standard error (MCSE) of the bridge sampling estimator, and (ii) introduce two complementary diagnostics–the Pareto-$\hat{k}$ tail index and a block-reshuffling bootstrap–to assess the adequacy of that MCSE estimate given finite MCMC sample size.
@@ 2 Bridge Sampling
One of the fastest and most reliable approaches to estimate the marginal likelihood is bridge sampling, first introduced by Bennett, (1976) and nicely reviewed by Gronau et al., (2017). The fundamental principle behind bridge sampling is the following identity
@@
$1=\frac{\int p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})h(\theta)g(\theta)}{\int p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})h(\theta)g(\theta)},$ (2)
which allows $p(\mathbf{y}\mid\mathcal{M})$ to be written as
@@
$p(\mathbf{y}\mid\mathcal{M})=\frac{\mathbb{E}_{g(\theta)}\left[p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})h(\theta)\right]}{\mathbb{E}_{\text{post}}\left[h(\theta)g(\theta)\right]},$ (3)
which consequently permits an approximation of the form
@@
$\hat{p}(\mathbf{y}\mid\mathcal{M})=\frac{\frac{1}{S_{2}}\sum_{i=1}^{S_{2}}p(\mathbf{y}\mid\tilde{\theta}_{i},\mathcal{M})p(\tilde{\theta}_{i}\mid\mathcal{M})h(\tilde{\theta}_{i})}{\frac{1}{S_{1}}\sum_{j=1}^{S_{1}}h(\theta^{*}_{j})g(\theta^{*}_{j})},$ (4)
where $S_{1}$ is the number of Markov chain Monte Carlo draws $\theta^{*}_{j}$ from the posterior distribution $p(\theta\mid\mathbf{y},\mathcal{M})$, $S_{2}$ is the number of draws $\tilde{\theta}_{i}$ from a proposal distribution $g(\theta)$ with known normalization constant, and $h(\theta)$ is the bridge function.
To ensure the effectiveness of bridge sampling, $g(\theta)$ has to closely approximate $p(\theta\mid\mathbf{y},\mathcal{M})$, be easy to sample from, and have a known normalization constant. For these reasons, it is typically chosen to be a multivariate normal distribution, denoted by $\operatorname{N}(\mu,\Sigma)$, where the mean vector $\mu$ and covariance matrix $\Sigma$ are estimated using additional posterior draws. Usually the original posterior draws are divided in two sets, with one set used as $\theta^{*}_{j}$ and the other set used to estimate the parameters of the proposal distribution $g(\theta)$. Instead of the normal distribution, other distributions with known normalization constants can be used, or the posterior may be warped to resemble more closely a normal distribution, and the marginal likelihood is then computed based on this warped distribution, which preserves the same normalizing constant as the original distribution.
The bridge function $h$ typically used is the one proposed by Meng and Wong, (1996):
$h(\theta)\propto\frac{1}{s_{1}p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})+s_{2}p(\mathbf{y}\mid\mathcal{M})g(\theta)},$ (5)
@@
Because (5) depends on the marginal likelihood $p(\mathbf{y}\mid\mathcal{M})$, the bridge sampling estimator is refined through an iterative scheme:
$\hat{p}^{(t+1)}(\mathbf{y}\mid\mathcal{M})=\frac{\dfrac{1}{S_{2}}\sum_{i=1}^{S_{2}}\dfrac{l_{2,i}}{s_{1}l_{2,i}+s_{2}\hat{p}^{(t)}(\mathbf{y}\mid\mathcal{M})}}{\dfrac{1}{S_{1}}\sum_{j=1}^{S_{1}}\dfrac{1}{s_{1}l_{1,j}+s_{2}\hat{p}^{(t)}(\mathbf{y}\mid\mathcal{M})}},\qquad\text{with}\;\hat{p}^{(0)}(\mathbf{y}\mid\mathcal{M})=1.$ (6)
where
$l_{1,j}\coloneqq\frac{p(\mathbf{y}\mid\theta^{*}_{j},\mathcal{M})p(\theta^{*}_{j}\mid\mathcal{M})}{g(\theta^{*}_{j})},$
and
$l_{2,i}\coloneqq\frac{p(\mathbf{y}\mid\tilde{\theta}_{i},\mathcal{M})p(\tilde{\theta}_{i}\mid\mathcal{M})}{g(\tilde{\theta}_{i})}.$
The iteration process continues until the magnitude of the update falls below a predefined threshold and, to prevent the algorithm from oscillating indefinitely, it is also set to terminate after a specified maximum number of iterations.
Unlike plain importance sampling, bridge sampling is less sensitive to the mismatch in tail thickness in proposal and posterior distribution. Specifically, both ratios in the numerator and denominator in Equation (6) are bounded by construction (see, e.g., Gronau et al., , 2017). The bounded ratios imply that both the numerator and denominator have finite variance, and thus the whole estimator has finite variance.
@@ 3 Monte Carlo Standard Error
As the first step to assess the accuracy of the bridge sampling estimate, we propose to compute Monte Carlo standard error (MCSE). The estimator $\hat{p}(\mathbf{y})$ given in Equation (4) is a ratio of two sample means:
$\hat{p}(\mathbf{y}\mid\mathcal{M})=\frac{\bar{N}}{\bar{D}},$
with
$\bar{N}$ $=\frac{1}{S_{2}}\sum_{i=1}^{S_{2}}N_{i},\quad N_{i}\,\coloneqq\,p(\mathbf{y}\mid\tilde{\theta}_{i},\mathcal{M})\,p(\tilde{\theta}_{i}\mid\mathcal{M})\,h(\tilde{\theta}_{i}),$
$\bar{D}$ $=\frac{1}{S_{1}}\sum_{j=1}^{S_{1}}D_{j},\quad D_{j}\,\coloneqq\,h(\theta^{*}_{j})\,g(\theta^{*}_{j}).$
To compute the variance of $\hat{p}(\mathbf{y}\mid\mathcal{M})$, we use the delta method, thus, the variance estimate is given by:
$\operatorname{Var}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)\approx\frac{1}{S_{1}}\left(\frac{\bar{N}}{\bar{D}}\right)^{2}\left(\frac{\operatorname{Var}(\bar{N})}{\bar{N}^{2}}+\frac{\operatorname{Var}(\bar{D})}{\bar{D}^{2}}-2\,\frac{\operatorname{Cov}(\bar{N},\bar{D})}{\bar{N}\,\bar{D}}\right),$ (7)
@@
$\operatorname{Var}(\bar{N})$ $=\frac{\operatorname{Var}(N_{i})}{S_{2}},$
$\operatorname{Var}(\bar{D})$ $=\frac{\operatorname{Var}(D_{j})}{S_{1}},$
$\operatorname{Cov}(\bar{N},\bar{D})$ $=\frac{\operatorname{Cov}(N_{i},D_{j})}{S_{1}}.$
The draws $\tilde{\theta}_{i}$ and $\theta^{*}_{j}$ are independent, and thus $\operatorname{Cov}(\bar{N},\bar{D})=0$. Substituting back, we get
$\operatorname{Var}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)\approx\frac{1}{S_{1}}\left(\frac{\bar{N}}{\bar{D}}\right)^{2}\left(\frac{\operatorname{Var}(N_{i})}{S_{2}\,\bar{N}^{2}}+\frac{\operatorname{Var}(D_{j})}{S_{1}\,\bar{D}^{2}}\right).$ (8)
If $\theta^{*}_{j}$ are drawn using MCMC, terms $D_{j}$ may also have significant autocorrelation and estimate for $\operatorname{Var}(\bar{D})$ should then take that into account. In the experiments we used MCMC effective sample size estimate by Vehtari et al., (2021). Finally
$\operatorname{MCSE}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)=\sqrt{\operatorname{Var}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)}.$ (9)
To compute the MCSE of the logarithm of the marginal likelihood estimate, $\log\hat{p}(\mathbf{y}\mid\mathcal{M})$, we use the property of the log-normal distribution to get
$\operatorname{Var}\left(\log\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)=\log\left(1+\frac{\operatorname{Var}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)}{\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)^{2}}\right).$
@@
$\operatorname{MCSE}\left(\log\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)=\sqrt{\operatorname{Var}\left(\log\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)}.$ (10)
The MCSE quantifies the uncertainty in the estimated (log) marginal likelihood due to the finite sample sizes $S_{1}$ and $S_{2}$, conditional on the proposal and the final bridge function when the iteration ends. In cases when the proposal is very different from the posterior or when the iterative algorithm fails to converge there can be additional variation. In Section 5, we present a block reshuffling approach which takes into account all the parts of the bridge sampling algorithm. However, based on the experiments reported in Section 6, the presented MCSE estimator alone is often sufficiently accurate.
4 Pareto-$\hat{k}$ Diagnostic
Let
$N_{i}\;=\;p(\mathbf{y}\mid\tilde{\theta}_{i},\mathcal{M})\,p(\tilde{\theta}_{i}\mid\mathcal{M})\,h(\tilde{\theta}_{i}),\qquad D_{j}\;=\;h(\theta^{*(j)})\,g(\theta^{*(j)}),$
so that the bridge estimator can be expressed as
$\hat{p}(\mathbf{y}\mid\mathcal{M})\;=\;\frac{\bar{N}}{\bar{D}},\qquad\bar{N}=\frac{1}{S_{2}}\sum_{i=1}^{S_{2}}N_{i},\qquad\bar{D}=\frac{1}{S_{1}}\sum_{j=1}^{S_{1}}D_{j}.$
Because $h(\cdot)$, $g(\cdot)$ and $p(\mathbf{y}\mid\cdot,\mathcal{M})\,p(\cdot\mid\mathcal{M})$ are all non-negative and bounded above on $\Theta$, there exist deterministic constants $B_{N},B_{D}<\infty$ such that
$0\;\leq\;N_{i}\leq B_{N},\qquad 0\;\leq\;D_{j}\leq B_{D}\quad\text{a.s.}$
Consequently, moments of distributions of $N_{i}$ and $D_{j}$ are all finite, and standard results (law of large numbers, delta method) guarantee that
$\sqrt{S_{1}}\!\left(\hat{p}(\mathbf{y}\mid\mathcal{M})-p(\mathbf{y}\mid\mathcal{M})\right)\;\xrightarrow{d}\;\operatorname{N}\bigl{(}0,\sigma^{2}\bigr{)},$
so, the variance-based Monte-Carlo standard error (MCSE) is asymptotically exact.
In practice the upper bounds $B_{N}$ and $B_{D}$ can be so large that, for realistic sample sizes $S_{1},S_{2}$, observations near either bound are exceedingly rare. The resulting finite sample behavior can be similar to that of heavy-tailed, or even infinite-variance, distributions: a single extreme realization may dominate $\bar{N}$ or $\bar{D}$ and inflate the estimator’s variance long before asymptotic behavior kicks in. Standard MCSE formulas, which assume finite variance, therefore give no early warning of this pre-asymptotic hazard.

### v2

@@ Abstract
In Bayesian statistics, the marginal likelihood is used for model selection and averaging, yet it is often challenging to compute accurately for complex models. Approaches such as bridge sampling, while effective, may suffer from issues of high variability of the estimates. We present how to estimate Monte Carlo standard error (MCSE) for bridge sampling, and how to diagnose the reliability of MCSE estimates using Pareto-$\widehat{k}$ and block reshuffling diagnostics without the need to repeatedly re-run full posterior inference. We demonstrate the behavior with increasingly more difficult simulated posteriors and many real posteriors from the posteriordb database.
Keywords: Marginal Likelihood, Normalizing Constant, Monte Carlo Standard Error
@@ 1 Introduction
Marginal likelihoods (model evidences) are often used for Bayesian model comparison (e.g. Bayes factors), model averaging, and calibrating prior and model components. At the same time, it is well understood that evidence-based workflows can be fragile in practice: marginal likelihood estimates can be extremely sensitive to prior specification and to numerical error, with different estimators’ behavior varying significantly across models and data. We do not advocate marginal likelihood as a universal model selection tool; rather, our goal is to reduce wasted computation and failure-by-silence by providing practical diagnostics for assessing the Monte Carlo reliability of evidence estimates computed from a finite number of posterior draws (see, e.g., Friel and Wyse,, 2012; Fourment et al.,, 2019 for broader reviews and cautions).
Let $\mathcal{M}$ denote a Bayesian model with parameter vector $\mathrm{\boldsymbol{\theta}}\in\Theta\subseteq\mathbb{R}^{p}$ (where $p\coloneqq\dim(\mathrm{\boldsymbol{\theta}})$) and observed data vector $\mathbf{y}=(y_{1},\dots,y_{n})^{\top}\in\mathbb{R}^{n}$.
$p(\mathbf{y}\mid\mathcal{M})=\int_{\Theta}p(\mathbf{y}\mid\mathrm{\boldsymbol{\theta}},\mathcal{M})\,p(\mathrm{\boldsymbol{\theta}}\mid\mathcal{M})\,d\mathrm{\boldsymbol{\theta}},$ (1)
which is generally intractable for realistic models and thus requires approximation. Among sampling-based methods, bridge sampling has a long history starting from the acceptance-ratio ideas of Bennett, (1976) and the general formulation of Meng and Wong, (1996), with connections to importance sampling and path sampling emphasized by Gelman and Meng, (1998). In modern Bayesian workflows bridge sampling has become a practical default, in part due to robust implementations and tutorials (e.g., Gronau et al.,, 2017; Gronau et al., 2020a, ).
Despite its practical appeal, bridge sampling is not “automatic” in the sense that a single run would always yield a numerically reliable estimate at typical posterior sample sizes; the estimator is a ratio of empirical averages whose stability is governed by overlap between the posterior and the proposal distribution and by the behavior of rare, extreme terms. In high-dimensional or strongly non-Gaussian posteriors, the distribution of the summands can exhibit severe pre-asymptotic behavior, so that a small number of atypical draws may dominate the estimate long before the asymptotic $\sqrt{S}$ regime becomes a useful approximation. Re-running full posterior inference to assess Monte Carlo variability is often computationally expensive, and practitioners therefore greatly benefit from diagnostics that can be computed from existing posterior draws and likelihood evaluations.
We develop a diagnostic workflow for bridge sampling that aims to answer a simple practical question: given the posterior draws already available, should we trust the reported marginal likelihood estimate and its Monte Carlo uncertainty? Concretely, we make three contributions:
  1. (i)
    We derive a closed-form Monte Carlo standard error (MCSE) estimator for the bridge sampling estimate (and its logarithm), accounting for autocorrelation in MCMC draws via effective sample size adjustments.
  2. (ii)
    We adapt Pareto tail-index diagnostics (Pareto-$\widehat{k}$) to the bridge sampling setting to flag finite-sample instability in the numerator and denominator contributions, complementing the MCSE.
  3. (iii)
    We introduce a block reshuffling diagnostic, a resampling procedure related to bootstrapping, demonstrating how this tool captures the skewness and non-normality of the estimator’s sampling distribution in finite samples, properties that the standard CLT-based MCSE cannot detect.
We demonstrate the behavior of these diagnostics using increasingly difficult simulated posteriors, and across a broad suite of real posteriors from the posteriordb database.
@@ 2 Bridge Sampling
Write the unnormalized posterior density as
@@
$\widetilde{\pi}(\mathrm{\boldsymbol{\theta}})\coloneqq p(\mathbf{y}\mid\mathrm{\boldsymbol{\theta}},\mathcal{M})\,p(\mathrm{\boldsymbol{\theta}}\mid\mathcal{M}),\qquad\pi(\mathrm{\boldsymbol{\theta}})\;=\;\frac{\widetilde{\pi}(\mathrm{\boldsymbol{\theta}})}{Z},\qquad Z\;\coloneqq\;p(\mathbf{y}\mid\mathcal{M})\;=\;\int_{\Theta}\widetilde{\pi}(\mathrm{\boldsymbol{\theta}})\,d\mathrm{\boldsymbol{\theta}}.$
Bridge sampling estimates the normalizing constant $Z$ by introducing a normalized proposal density $g(\mathrm{\boldsymbol{\theta}})$ (from which we can both sample and evaluate pointwise) whose support contains that of the posterior $\pi(\mathrm{\boldsymbol{\theta}})$, together with a bridge function $h(\mathrm{\boldsymbol{\theta}})$ (also denoted $\omega(\mathrm{\boldsymbol{\theta}})$ in parts of the literature). Formally, $h(\mathrm{\boldsymbol{\theta}})$ is a user-chosen auxiliary function defined on the common support of $\pi$ and $g$ such that
$0<\left|\int_{\Theta}h(\mathrm{\boldsymbol{\theta}})\pi(\mathrm{\boldsymbol{\theta}})g(\mathrm{\boldsymbol{\theta}})\,d\mathrm{\boldsymbol{\theta}}\right|<\infty.$
Provided this condition holds, $h(\mathrm{\boldsymbol{\theta}})$ acts as an algebraic device satisfying the exact identity
@@
$Z\;=\;\frac{\mathbb{E}_{g}\!\left[\widetilde{\pi}(\mathrm{\boldsymbol{\theta}})\,h(\mathrm{\boldsymbol{\theta}})\right]}{\mathbb{E}_{\pi}\!\left[g(\mathrm{\boldsymbol{\theta}})\,h(\mathrm{\boldsymbol{\theta}})\right]},$ (2)
where
$\mathbb{E}_{g}[f(\mathrm{\boldsymbol{\theta}})]\coloneqq\int_{\Theta}f(\mathrm{\boldsymbol{\theta}})\,g(\mathrm{\boldsymbol{\theta}})\,d\mathrm{\boldsymbol{\theta}}$
and
$\mathbb{E}_{\pi}[f(\mathrm{\boldsymbol{\theta}})]\coloneqq\int_{\Theta}f(\mathrm{\boldsymbol{\theta}})\,\pi(\mathrm{\boldsymbol{\theta}})\,d\mathrm{\boldsymbol{\theta}}.$
Hence, $h$ defines the resulting Monte Carlo estimator and strongly affects its variance through how it reweights regions where $\pi$ and $g$ overlap, with many familiar estimators arising as special cases of bridge sampling via particular choices of $h$ (e.g., importance sampling and reciprocal-importance/harmonic-mean–type estimators).
Using the identity (2), we obtain the empirical estimator
@@
$\widehat{Z}\;=\;\widehat{p}(\mathbf{y}\mid\mathcal{M})\;=\;\frac{\frac{1}{S_{2}}\sum_{i=1}^{S_{2}}\widetilde{\pi}(\widetilde{\mathrm{\boldsymbol{\theta}}}_{i})\,h(\widetilde{\mathrm{\boldsymbol{\theta}}}_{i})}{\frac{1}{S_{1}}\sum_{j=1}^{S_{1}}h(\mathrm{\boldsymbol{\theta}}_{j}^{*})\,g(\mathrm{\boldsymbol{\theta}}_{j}^{*})},$ (3)
where $\mathrm{\boldsymbol{\theta}}_{j}^{*}\sim\pi(\mathrm{\boldsymbol{\theta}})$ are (typically MCMC) $S_{1}$ posterior draws and $\widetilde{\mathrm{\boldsymbol{\theta}}}_{i}\sim g(\mathrm{\boldsymbol{\theta}})$ are $S_{2}$ independent proposal draws.
Figure 1: Intuition of the bridge sampling estimator. It relies on the ratio of the unnormalized posterior $\widetilde{\pi}(\mathrm{\boldsymbol{\theta}})$ (solid line) and the proposal density $g(\mathrm{\boldsymbol{\theta}})$ (dashed line), and its stability depends on the overlap (shaded region) between the two distributions.
The practical performance of bridge sampling is governed by overlap between $\pi(\mathrm{\boldsymbol{\theta}})$ and $g(\mathrm{\boldsymbol{\theta}})$ (see Figure 1), which is frequently chosen as a multivariate normal distribution in the unconstrained parameter space, with mean and covariance estimated from posterior draws (see, e.g., Gronau et al.,, 2017; Gronau et al., 2020a, for the standard workflow). Because proposal fitting from posterior draws can introduce bias and/or lead to over-optimistic stability assessments if the same draws are reused, implementations often split posterior draws into subsets (one subset to fit $g$, one subset to estimate $Z$), trading reduced bias for increased Monte Carlo variability (Overstall and Forster,, 2010; Wong et al.,, 2020). For a comprehensive tutorial and step-by-step implementation details, we refer the reader to Gronau et al., (2017).
As some models may present strongly non-Gaussian or multimodal posteriors, leading to poor overlap with normal proposals, a complementary line of work focuses on improving overlap by transforming (warping) the target to match more closely a tractable reference distribution (see, e.g., warp bridge sampling (Meng and Schilling,, 2002; Gronau et al., 2020b, ) and Warp-U–style transformations for multimodal targets (Wang et al.,, 2022)). Our focus is not to introduce new proposals or warps, but rather to provide diagnostics that indicate when the default workflow is already reliable and when additional computational investment (more posterior draws, better proposals/warps, or alternative evidence estimators) is warranted.
A widely used (near-optimal) choice of $h(\mathrm{\boldsymbol{\theta}})$ is the one proposed by Meng and Wong, (1996),
$h(\mathrm{\boldsymbol{\theta}})\propto\frac{1}{s_{1}\widetilde{\pi}(\mathrm{\boldsymbol{\theta}})+s_{2}\,Z\,g(\mathrm{\boldsymbol{\theta}})},$ (4)
@@
More generally, while the asymptotically optimal bridge function under independence is proportional to the reciprocal of a mixture of the two (unnormalized) densities (Bennett,, 1976; Meng and Wong,, 1996), because this optimal form depends on the unknown $Z$, it is implemented in practice via the fixed-point iteration in (5).
Defining
$l_{1,j}\coloneqq\frac{\widetilde{\pi}(\mathrm{\boldsymbol{\theta}}^{*}_{j})}{g(\mathrm{\boldsymbol{\theta}}^{*}_{j})},\qquad l_{2,i}\coloneqq\frac{\widetilde{\pi}(\widetilde{\mathrm{\boldsymbol{\theta}}}_{i})}{g(\widetilde{\mathrm{\boldsymbol{\theta}}}_{i})},$
the standard update takes the form
$\widehat{Z}^{(t+1)}=\frac{\dfrac{1}{S_{2}}\sum_{i=1}^{S_{2}}\dfrac{l_{2,i}}{s_{1}l_{2,i}+s_{2}\widehat{Z}^{(t)}}}{\dfrac{1}{S_{1}}\sum_{j=1}^{S_{1}}\dfrac{1}{s_{1}l_{1,j}+s_{2}\widehat{Z}^{(t)}}},\qquad\text{with}\;\widehat{Z}^{(0)}=1,$ (5)
iterating until the update is smaller than a tolerance (or until a maximum number of iterations is reached).
Bridge sampling is typically less sensitive than plain importance sampling to modest tail mismatch because the update in (5) involves bounded ratios (see, e.g., Gronau et al.,, 2017). However, boundedness alone does not guarantee stable finite-sample behavior: when overlap is poor, the effective range of the summands can be so wide that rare extreme realizations dominate at practical sample sizes $S_{1}$ and $S_{2}$. The remainder develops MCSE estimation and complementary diagnostics aimed at detecting exactly these pre-asymptotic failure modes without requiring repeated posterior inference.
@@ 3 Monte Carlo Standard Error
As the first step to assess the precision of the bridge sampling estimate, we propose to compute Monte Carlo standard error (MCSE). The estimator $\widehat{Z}$ given in Equation (3) is a ratio of two sample means: $\widehat{Z}=\bar{N}/\bar{D}$, with
$\bar{N}$ $=\frac{1}{S_{2}}\sum_{i=1}^{S_{2}}N_{i},\quad\text{where}\quad N_{i}\,\coloneqq\,\widetilde{\pi}(\widetilde{\mathrm{\boldsymbol{\theta}}}_{i})h(\widetilde{\mathrm{\boldsymbol{\theta}}}_{i}),$ (6)
$\bar{D}$ $=\frac{1}{S_{1}}\sum_{j=1}^{S_{1}}D_{j},\quad\text{where}\quad D_{j}\,\coloneqq\,h(\mathrm{\boldsymbol{\theta}}^{*}_{j})\,g(\mathrm{\boldsymbol{\theta}}^{*}_{j}).$
To compute the variance of $\widehat{Z}$, we use the delta method, thus, the variance estimate is:
$\operatorname{Var}\left(\widehat{Z}\right)\approx\frac{1}{S_{1}}\left(\frac{\bar{N}}{\bar{D}}\right)^{2}\left(\frac{\operatorname{Var}(\bar{N})}{\bar{N}^{2}}+\frac{\operatorname{Var}(\bar{D})}{\bar{D}^{2}}-2\,\frac{\operatorname{Cov}(\bar{N},\bar{D})}{\bar{N}\,\bar{D}}\right),$ (7)
@@
$\operatorname{Var}(\bar{N})\coloneqq\frac{\operatorname{Var}(N_{i})}{S_{2}},\qquad\operatorname{Var}(\bar{D})\coloneqq\frac{\operatorname{Var}(D_{j})}{S_{1}},\qquad\operatorname{Cov}(\bar{N},\bar{D})\coloneqq\frac{\operatorname{Cov}(N_{i},D_{j})}{S_{1}}.$
The draws $\widetilde{\mathrm{\boldsymbol{\theta}}}_{i}$ and $\mathrm{\boldsymbol{\theta}}^{*}_{j}$ are independent, and thus $\operatorname{Cov}(\bar{N},\bar{D})=0$. Substituting back, we get
$\operatorname{Var}\left(\widehat{Z}\right)\approx\left(\frac{\bar{N}}{\bar{D}}\right)^{2}\left(\frac{\operatorname{Var}(N_{i})}{S_{2}\,\bar{N}^{2}}+\frac{\operatorname{Var}(D_{j})}{S_{1}\,\bar{D}^{2}}\right).$ (8)
If $\mathrm{\boldsymbol{\theta}}^{*}_{j}$ are drawn using MCMC, terms $D_{j}$ may also have significant autocorrelation and estimate for $\operatorname{Var}(\bar{D})$ should then take that into account. In the experiments we used MCMC effective sample size estimate by Vehtari et al., (2021). Finally
$\operatorname{MCSE}\left(\widehat{Z}\right)=\sqrt{\operatorname{Var}\left(\widehat{Z}\right)}$. (9)
To compute the MCSE of the logarithm of the marginal likelihood estimate, $\log\widehat{Z}$, we use the property of the log-normal distribution to get
$\operatorname{Var}\left(\log\widehat{Z}\right)=\log\left(1+\frac{\operatorname{Var}\left(\widehat{Z}\right)}{\widehat{Z}^{2}}\right).$
@@
$\operatorname{MCSE}\left(\log\widehat{Z}\right)=\sqrt{\operatorname{Var}\left(\log\widehat{Z}\right)}$. (10)
The MCSE quantifies the uncertainty in the estimated (log) marginal likelihood due to the finite number of draws $S_{1}$ and $S_{2}$, conditional on the proposal and the final bridge function when the iteration ends. In cases when the proposal is very different from the posterior or when the iterative algorithm fails to converge, there can be additional variation. In Section 5, we present a block reshuffling approach which takes into account all the parts of the bridge sampling algorithm. However, based on the experiments reported in Section 6, the presented MCSE estimator alone is often sufficiently precise.
4 Pareto-$\widehat{k}$ Diagnostic and Pre-asymptotic Stability